OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x – 3y = –3 2x – 6y = –6 –1 0 1 2

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (zzr0ck3r):

5x – 3y = –3 2x – 6y = –6 multiply the first equation by -2 then add the two equations -10x+6x = 6 2x - 6y = -6 add them -8x = 0 x = 0 so -3y = -3 and -6y = -6 when y = ?

OpenStudy (anonymous):

1

OpenStudy (zzr0ck3r):

yes